OpenStudy (lulubj):

Find the product. Write your answer in exponential form. 1^2 * 1^2

OpenStudy (lulubj):

would the answer just be 1?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

yes 1

OpenStudy (welshfella):

Here the exponent will be 2+2 so its 1^ what? 1^4 = 1 yes

OpenStudy (welshfella):

because 1*1*1*1 = 1

OpenStudy (lulubj):

i put 1 and they marked it wrong :(

sam (.sam.):

\[1^2*1^2=1^{2+2}=1^4\]

OpenStudy (welshfella):

hmmm yes they wanted it in exponent form has sam has done

OpenStudy (lulubj):

ohh damnn

OpenStudy (lulubj):

Well thanks guys

OpenStudy (misty1212):

maybe there is a typo somewhere there are you sure it is \[1^2\times 1^2\] just like that?

OpenStudy (welshfella):

maybe its the letter I and not the number 1

OpenStudy (misty1212):

could it be \[10^2\times 10^2\] or \[x^2\times x^2\] or sommat?

sam (.sam.):

The answer was \(1^4\) not 1

OpenStudy (misty1212):

oh yeah or \[l^2\times l^2\]

OpenStudy (misty1212):

\[1^4\] is no more in "exponential form" than \(1^0=1^1=1^2=1^3=...\)